# Communicating the gravitational-wave discoveries of the LIGO–Virgo–KAGRA Collaboration

### v2

@@ Abstract
The LIGO–Virgo–KAGRA (LVK) Collaboration has made breakthrough discoveries in gravitational-wave astronomy, a new field of astronomy that provides a different means of observing our Universe. Gravitational-wave discoveries are possible thanks to the work of thousands of people from across the globe working together. In this article, we discuss the range of engagement activities used to communicate LVK gravitational-wave discoveries and the stories of the people behind the science using the activities surrounding the release of the third Gravitational Wave Transient Catalog as a case study.
@@ 1 Introduction: Context and Objectives
In 2015, one hundred years after Einstein first calculated the properties of GWs in his theory of general relativity, the first direct observation of GWs was made [1]. This discovery was confirmation for Einstein’s theory, an experimental triumph, and the beginning of a new era for astronomy. The signal GW150914 came from two black holes, each about $30$ times the mass of our Sun, coalescing more than a billion light-years away. This was the first time such a binary black hole system was found, the first time black holes of this size were discovered, and the first time two black holes were observed to inspiral and merge. The discovery was a major news event around the world [2], building on the public’s enduring fascination with Einstein and black holes.
@@
Since 2015, the LIGO–Virgo–KAGRA (LVK) Collaboration, which operates the international network of GW observatories (the two LIGO sites in the US, Virgo in Italy, KAGRA in Japan and GEO 600 in Germany) has made many further GW discoveries. The field has grown rapidly, with the third GW Transient Catalog (GWTC-3) increasing the number of probable detections to $90$ [3]. The latest observing run is ongoing. These observations have revealed a diverse range of black hole and neutron star binaries, and revolutionised our understanding of the astrophysics of these sources.
@@
Surrounding each results release are numerous engagement and communication activities. These share the wonder of discovery, what we can learn with GWs, and the technological advances that enable these groundbreaking measurements. Engagement and communication activities are carried out through the effort and enthusiasm of individual scientists, collaboration-wide teams, and partnerships between scientists and non-scientists, all of which complements the work of professional science communicators at the observatories. Matching the global composition of the LVK, selected resources have been translated into over $20$ languages.
@@
Discoveries also provide an opportunity to focus on the people behind the science. The LVK includes around $2000$ people from diverse backgrounds across the globe; from designing instrumentation to operating the kilometre-scale laser interferometers, from coding analysis algorithms to calculating GW emission. By making connections with the people within the LVK, we show the human side of how science works, and provide possible role models for young people interested in science and technology.
@@
In this article, we detail communication methods used by the LVK. We review the range of engagement activities used to communicate LVK GW discoveries and the stories of the people behind the science, with specific examples for GWTC-3.
@@ 2 General communication methods
2.1 The people behind the science
Figure 1: Frequency of Humans of LIGO participants’ institution (top) and hometown (bottom) county. Scientists come from more diverse locations than just the institutions of the LVK.
@@ 2.1 The people behind the science
Discovery science is not just about the results, it is also about the people behind them. Highlighting individuals within the LVK can provide STEM role models, challenge stereotypes, and help to foster a welcoming environment.
Figure 2: Antimatter comic showing LIGO scientists performing detector checks (credit: Nutsinee Kijbunchoo).
The LIGO Magazine (www.ligo.org/magazine) is published twice annually. Content is commissioned and edited by volunteers from the LIGO Scientific Collaboration and it has a professional, high-quality design for print and online PDF. Since the first edition in 2012, the LIGO Magazine has featured $470$ authors, representing $237$ workplaces in $23$ countries. Authors are from the GW community and beyond, and the target audience is both collaboration members and the interested public. Articles include perspectives of working on GW science, personal stories, opinion pieces, and advice columns.
@@
Humans of LIGO (humansofligo.blogspot.com) showcases the lives, backgrounds, and inspirations of LVK members and GW scientists through blog posts. Each post consists of an image of the participant, a direct quote about their experiences, and a short biography including job description and hobbies. Since its inception in 2018, the blog has featured $59$ participants from $11$ different counties, working at $42$ institutions (Figure 1). The blog posts are shared on LIGO social media, and have reached over $53,000$ views (as of May 2024).
@@
Antimatter web comics (antimatterwebcomics.com) by Nutsinee Kijbunchoo features day-to-day experiences and current affairs from the GW community and beyond (Figure 2). Mental health and graduate-student life are focuses of many comics. Comics have been regularly featured in the LIGO Magazine.
@@
The LIGO-India blog, Glorious Women [4] showcases STEM role models, highlighting women from across the LVK. Behind-the-scenes interviews feature the daily lives of students, their motivation, contributions and challenges [5].
2.2 Academic audiences
@@
Communication with academics is primarily through scholarly articles. Since GW astronomy is a new field, the LVK have written dedicated articles on introductory GW data analysis [6], the structure of our data releases [7, 8], and the basic physics of a binary black hole coalescence (suitable for undergraduate teaching) [9].
@@
Data releases are coordinated by the GW Open Science Center (GWOSC), which hosts raw data, links to LVK publications and further data products (gwosc.org). GWOSC maintains lists of open-source analysis software and provides teaching on data analysis through annual Open Data Workshops. The workshops are delivered in a hybrid format; with 200200200200300300300300 people participating in-person each year, and over $7700$ people enrolling online over the last four years. Workshop materials are openly available online.
To accompany the release of key papers, the LVK organises webinars through Zoom, with attendees able to submit questions. Recordings are uploaded to YouTube.
@@
The LVK also engages with the academic community to promote proper recognition of its three constituent collaborations. There is currently a bias in scholarly literature: an abbreviated narrative attributes credit to LIGO, excluding Virgo and KAGRA from LVK achievements, with consequences for scientific careers, accuracy of public media and interactions with funding agencies. A year-long project [10] studied this cognitive bias, and tried to educate the community.
2.3 Formal classroom education
@@
LVK members have maintained long-term efforts to introduce GWs and related topics into formal classroom education. Examples include an Educator’s Guide [11], courses for community college teachers, high school resources, and undergraduate lab demonstrations [12].
Einstein-First (www.einsteinian physics.com) has developed an eight-year science curriculum that introduces quantum concepts and general relativity at primary school, followed by GWs, black holes, cosmology and quantum measurement in middle school [13]. Efforts are underway to introduce Einstein-First in Australia, India, USA and Brazil.
2.4 Science Summaries and press releases
@@
LVK Science Summaries are a long-standing effort to explain LVK publications at a technical level typically higher than in the popular press but still accessible for enthusiastic lay readers. They are published on www.ligo.org, distributed as print copies at outreach events, and shared with journalists. Summaries are translated into a variety of languages by LVK volunteers, with discovery papers often receiving the most attention (e.g., the GW190521 summary was translated into 16 languages) [14].
@@
Press releases to accompany discovery announcements are prepared centrally, including quotations from a selection of LVK members, and then adapted by member institutions (often including quotations from their scientists). These are frequently drawn upon for news articles.
2.5 In-person interactions: Science festivals, museums and visitor centers
@@
In-person outreach to the general public reaches fewer people than other means, but it can yield impactful interactions, and may be perceived as having more value [15].
@@
LVK members participate in science festivals, classroom visits, the creation of science exhibits (such as interactive detector models [16, 12], visitor centers and tours of the GW observatories. The face-to-face format allows staff to customize the experience to the visitor.
@@
LVK’s visitor centers merge artifacts from the detectors with interactive exhibits that put the visitor in the role of a scientist or engineer. The exhibits explore the underlying science and engineering of GW detectors, from the importance of pendula to the concepts underpinning gravity. In-person and virtual tours of the detectors give a glimpse to the heart of operations. Over $330,000$ visitors have interacted with the LVK visitor centers; 2024 attendance is projected to reach $30,000$.
Science festivals and local outreach efforts around the world connect with classrooms and the general public. Outreach is performed by seasoned scientists, engineers, graduate students and undergraduates who have only recently learned about GWs. Such outreach efforts positively impact the undergraduates who perform the outreach [17, 18, 19].
2.6 Graphics
@@
Figure 3: The detector data for GW150914, and an artistic representation of its binary black hole source (credit: LIGO/National Science Foundation/Aurore Simmonet/Somona State University). The source is shown at three stages, corresponding to the portions of the signal beneath.
@@
Unlike electromagnetic observatories, GW data does not naturally lend itself to the stunning images that are commonplace throughout astronomy. Nevertheless, visualisations, infographics and cartoons [20]form key parts of our outreach activities, e.g., infographics displaying key facts are produced to accompany discoveries; simplified plots of results may be created for Science Summaries, and artistic representations of our sources may accompany press releases.
@@
Videos and images are produced from numerical-relativity simulations of sources and the GWs they emit. Since new simulations are often produced to better understand the physics of novel sources, fresh visualisations have often accompanied discovery announcements.
@@
Images accompanying our discovery announcements have been selected as NASA Astronomy Pictures of the Day, e.g., to accompany our first detection, an artistic representation of a binary black hole coalescence (Figure 3) and a corresponding numerical-relativity simulation were featured (11 and 12 February 2016), and to accompany GW190521, an artistic representation of the GW emission from two spinning black holes was featured (8 September 2020).
2.7 Multisensory resources: sound and touch

### v3

@@ Abstract
The LIGO–Virgo–KAGRA (LVK) Collaboration has made breakthrough discoveries in gravitational-wave astronomy, a new field that provides a different means of observing our Universe. Gravitational-wave discoveries are possible thanks to the work of thousands of people from across the globe working together. In this article, we discuss the range of engagement activities used to communicate LVK gravitational-wave discoveries and the stories of the people behind the science, using the activities surrounding the release of the third Gravitational-Wave Transient Catalog as a case study.
@@ 1 Introduction: Context and Objectives
In 2015, $100$ years after Einstein first calculated the properties of GWs in his theory of general relativity, the first direct observation of GWs was made [Abbott et al., 2016b]. This discovery was confirmation for Einstein’s theory, an experimental triumph, and the beginning of a new era for astronomy. The signal GW150914 came from two black holes, each about $30$ times the mass of our Sun, coalescing more than a billion light-years away. This was the first time such a binary black hole system was found, the first time black holes of this size were discovered, and the first time two black holes were observed to inspiral and merge. The discovery was a major global news event [Key et al., 2016], building on the public’s fascination with Einstein and black holes.
@@
Since 2015, the LIGO–Virgo–KAGRA (LVK) Collaboration, which operates the international network of ground-based laser-interferometric GW observatories (the two LIGO sites in the US [Aasi et al., 2015], Virgo in Italy [Acernese et al., 2015], KAGRA in Japan [Akutsu et al., 2021] and GEO 600 in Germany [Dooley et al., 2016] has made many further GW discoveries. The field has grown rapidly, with the third GW Transient Catalog (GWTC-3) increasing the number of probable detections to $90$ [Abbott et al., 2023a]. The latest observing run is ongoing. These observations have revealed a diverse range of black hole and neutron star binaries, and revolutionised our understanding of these sources’ astrophysics.
@@
Surrounding each results release are numerous engagement and communication activities. These share the wonder of discovery, what can be learned from GWs, and the technological advances that enable these groundbreaking measurements. Matching the global composition of the LVK, selected resources have been translated into over $20$ languages.
Engagement, communication and education activities are carried out by a wide variety of individuals: from those who see themselves as scientists and engineers with a passion for outreach, to those who view themselves as outreach experts with a passion for science. These individuals make commitments as LVK members, and join together in collaboration-wide teams and outreach partnerships.
@@
Discoveries also provide an opportunity to focus on the people behind the science. The LVK includes around $2000$ people from diverse backgrounds across the globe. Their activities range from designing instrumentation to operating the kilometre-scale laser interferometers, from coding analysis algorithms to calculating GW emission. Making connections with the people within the LVK reveals the human side of how science works, and provides possible role models for young people interested in science, technology, engineering and mathematics (STEM).
Communicating LVK discoveries faces several challenges. Some, such as explaining complicated ideas using non-technical language or engaging with hard-to-reach communities, are common to other science communications. However, LVK science also faces specific challenges including explaining the concept of GWs, which are alien to everyday experiences; constructing a narrative around the work of a large collaboration, where it is often not possible to identify distinct contributions of individuals, and maintaining interest as the field develops from an era of first detections to using large catalogs to make statistical statements. Each of these challenges also provides an opportunity, enabling us to increase the scientific literacy of those reached: explaining the results of LVK work, how they are produced, and how modern science makes discoveries through carefully combining many pieces of research.
@@
In this article, we detail communication methods used by the LVK. We first review the range of engagement activities used to communicate LVK GW discoveries and the stories of the people behind the science, and then provide specific examples of activities used for GWTC-3. We conclude with reflections upon these activities and the challenges of communicating LVK discoveries.
@@ 2 General communication methods
Table 1: A selection of LVK communication activities and examples we discuss in this article.
Activity Examples Audiences Media
Sharing the people behind the science LIGO Magazine; Humans of LIGO; Antimatter comics; LIGO-India blog LVK members; other academics; undergraduate students; school students; general public Print; online writing; graphics
Engaging academia Journal articles; Open Data Workshops; data releases; webinars; direct interventions LVK members; other academics; undergraduate students; interested public Academic resources; live and recorded talks; online course materials; online documentation; online correspondence
Supporting formal classroom education Educator’s guide; Einstein-First; Space Public Outreach Team School educators; undergraduate students; school children Online written resources; face-to-face communication
Writing reference texts Science Summaries; press releases; news items Undergraduate students; science journalists; interested public Online writing; print
Interacting in person Science festivals; museum exhibits; visitor centres School children; general public Face-to-face communication; talks; live demonstrations; interactive exhibits
Creating graphics Infographics; simulation visualisations; artistic impressions; Masses in the Stellar Graveyard; LVK Orrery Academics; undergraduate students; journalists; general public Images; videos
Employing multisensory resources Sounds of Spacetime; Tactile Universe; Low Mass Beats General public Audio; 3D-printed materials
Combining art and music with science GravitySynth; GWSciArt; Celebrating Einstein festival General public Visual arts; music
Communicating through social and non-traditional media Podcasts; social-media posting; Reddit Ask-me-anything Science journalists; interested public; general public Online correspondence; online writing; audio; graphics; videos
Using interactive technologies Laser Labs games and apps; Black Hole Hunter; Mission Gravity School children; general public Interactive software
@@
The LVK is a consortium of its three constituent collaborations, each with its own internal organisational structure. Communications efforts are primarily led by a division-level coordinator or chair, who oversees various working groups, committees, and individual efforts working towards shared objectives. Activities organised through these groups are typically undertaken by teams of individuals, who then report their progress through regular meetings, LVK-wide conferences, and annual reports. Some activities, like coordination of press releases, may be done by dedicated communications teams in each of the three collaborations, but most activities are done by volunteers from the LVK. Division leads are responsible for ensuring that these activities complement each other and align with parallel efforts across the observatory sites.
The LVK consists of many individuals with different backgrounds and job roles working in institutions around the world and funded by different sources. This results in a diverse set of public-engagement activities run by different teams. While these may share a broad goal of communicating GW science, they may differ in specific goals, e.g., encouraging schoolchildren to study STEM or informing science journalists. The different activities complement each other, and provide a breadth of resources and expertise for LVK members to draw upon. Below we group in terms of activities (Table 1), reflecting how these are organised within the LVK by different teams.
2.1 Sharing the people behind the science
@@
Discovery science is not just about the results, it is also about the people behind them. Highlighting individuals within the LVK can provide STEM role models, challenge stereotypes, and help to foster a welcoming environment by promoting diversity and inclusion.
The LIGO Magazine (www.ligo.org/magazine), published twice annually, is commissioned and edited by a team of volunteers from the LIGO Scientific Collaboration. It was conceived to build connections across the diverse GW community and inform the interested public. Since the first edition in 2012, the Magazine has featured 470 authors, representing $237$ workplaces in $23$ countries (as of Issue 24, March 2024). Articles include perspectives of working on GW science, personal stories, opinion pieces, science explainers, and advice columns. The Magazine provides opportunities for early-career scientists to discuss their behind-the-scenes experiences of working on big GW results, as well as an avenue for current and former LVK members to share their perspectives of working in academia and beyond. The LIGO Magazine has a professional, high-quality design. For each edition, $500$ hard copies are printed. Approximately half are sent to LVK Collaboration meetings for participants to take back to their home institutions; the rest are distributed between detector sites and other LVK institutions worldwide. Print copies are also used within LVK as handouts for government representatives, students, and to visitors at detector sites and institutions. The full archive is also online as PDF downloads.
Figure 1: Frequency of Humans of LIGO participants’ institution (top) and hometown (bottom) county.
@@
Humans of LIGO (humansofligo.blogspot.com) showcases the lives, backgrounds, and inspirations of LVK members and GW scientists through blog posts. Each post consists of an image of the participant, a direct quote about their experiences, and a short biography including job description and hobbies. The blog posts are shared on LIGO social media, and have reached over $53,000$ views (as of May 2024). Since its inception in 2018, the blog has featured $59$ participants from $11$ different countries, working at $42$ institutions (Figure 1). While profiled scientists come from more diverse locations than just those of LVK institutions, they are still concentrated, reflecting the inequitable distribution of scientific research [King, 2004]. A future goal is to add profiles for scientists from other backgrounds and demonstrate potential opportunities within STEM for people across the world.
Figure 2: Antimatter comic showing LIGO scientists performing detector checks (credit: Nutsinee Kijbunchoo). A version of this comic featured in the CQG+ blog accompanying a paper describing detector characterization for GW150914 [Abbott et al., 2016a].
@@
Antimatter web comics (antimatterwebcomics.com) by Nutsinee Kijbunchoo features day-to-day experiences and current affairs from the GW community and beyond (Figure 2). Mental health and graduate-student life are focuses of many comics. Based upon webpage views (August 2024), the most popular two comics are about the discovery of the first binary neutron star signal and about living with depression, demonstrating that both communicating discoveries and sharing the lives of the scientists behind them have appeal. Comics have been regularly featured in the LIGO Magazine.
@@
The LIGO-India blog, Glorious Women [LIGO-India, 2024a] showcases STEM role models, highlighting women from across the LVK. Behind-the-scenes interviews feature the daily lives of students, their motivation, contributions and challenges [LIGO-India, 2024b].
2.2 Engaging academia
@@
Communication with academics is primarily through scholarly articles. Since GW astronomy is new, the LVK have written dedicated articles on introductory GW data analysis [Abbott et al., 2020a], our data releases [Abbott et al., 2021, 2023b], and the basic physics of a binary coalescences (suitable for undergraduate teaching) [Abbott et al., 2017].
@@
Data releases are coordinated by the GW Open Science Center (GWOSC), which hosts raw data, links to LVK publications and further data products (gwosc.org). GWOSC maintains lists of open-source analysis software, and provides teaching on data analysis through annual Open Data Workshops. Workshops are delivered in a hybrid format, with 200200200200300300300300 people participating in-person each year, and over $7,700$ people enrolling online over the last four years. Workshop materials are openly available online.
To accompany the release of key papers, the LVK organises Zoom webinars, with attendees able to submit questions. These provide convenient, direct communication with other scientists, and reach a larger audience than many in-person conferences. Recordings are uploaded to YouTube. YouTube views range from $400$ to $3,400$ (August 2024), with the older recordings accumulating more views.
@@
The LVK also engages with the academic community to promote proper recognition of its three constituent collaborations. There is currently a bias in scholarly literature: an abbreviated narrative attributes credit to LIGO, excluding Virgo and KAGRA from LVK achievements, with consequences for scientific careers, accuracy of media and interactions with funding agencies. A year-long project [Barneo et al., 2024] studied this cognitive bias, and tried to educate the community. While intervention has encouraged authors to include proper attribution in their work, it has yet to be seen if this has a long-term impact of reducing incomplete attribution.
2.3 Supporting formal classroom education
@@
LVK members have maintained long-term efforts to introduce GWs and related topics into formal classroom education. Examples include an Educator’s Guide [Edeon STEM Learning, 2016], courses for community-college teachers, high-school resources, and undergraduate lab demonstrations [Gardner et al., 2022].
Einstein-First (www.einsteinian physics.com) has developed an eight-year spiral curriculum Eight Steps to Einstein’s Universe for students aged 77778888 to 1515151516161616 years old. This introduces fundamental concepts (e.g., atoms, molecules, photons and phonons), and in the last year of compulsory science education, a module about GWs, black holes, climate change and Hubble’s law [Popkova et al., 2023]. The Einstein-First team is upskilling almost $100$ teachers through micro-credentials courses with many teachers now delivering the program. Micro-credential training successfully empowered teachers: $96\%$ agreed or strongly agreed both that they understood why Einsteinian concepts were important to the school curriculum, and that they would feel confident running Einstein-First activities for their students [Kaur et al., 2023a, b]. Efforts are underway to introduce Einstein-First in Australia, India, USA, Greece, and Brazil.
Outreach to schools and STEM education programs provides an avenue for sharing GW discoveries, engaging students and teachers, and additionally provides science-communication training for GW researchers. Assessment of the impact of the Space Public Outreach Team (SPOT) program shows positive outcomes for students at all levels for SPOT programs across a range of geographic locations, benefitting the college-student presenters with improved scientific knowledge and presentation skills, and the audience of school children with greater engagement with science [Key et al., 2024; Des Jardins et al., 2020].
2.4 Writing reference texts
Figure 3: Number of Science Summaries by year of posting, and their number of translations. There were fewer papers in 2021–2024 while waiting for the results of the latest observing run.
LVK Science Summaries and press releases are both Collaboration-coordinated activities to provide written summaries of our discoveries.
@@ 2.4 Writing reference texts
LVK Science Summaries are a long-standing effort to explain LVK publications at a technical level typically higher than in the popular press but still accessible for enthusiastic lay readers. They are published on www.ligo.org, distributed as print copies at outreach events, and shared with journalists. Summaries for discovery announcements may get thousands to tens of thousands of views around publication. Summaries are translated into a variety of languages by LVK volunteers, with discovery papers often receiving the most attention (e.g., the GW190521 summary was translated into 16 languages) [Keitel et al., 2021]. Numbers of summaries and translations over time are displayed in Figure 3, demonstrating significant growth in translation activity especially during the LVK’s third observing run (2019–2020). However, translation activity is typically less diverse for summaries of lower-profile (non-discovery) papers, demonstrating a similar pattern of enthusiasm by LVK translators and the target audience
@@
Press releases to accompany discovery announcements are prepared centrally, including quotations from a selection of LVK members, and then adapted by member institutions (often including quotations from their scientists). As a global collaboration, it is not possible to find a single time for press releases to be made public that works for all time zones, which can make it challenging to coordinate local coverage.
Both Science Summaries and press releases have been drawn upon by journalists [e.g., Carlise, 2021]. Translation of Science Summaries is especially important for use in non-English media.
2.5 Interacting in person
@@
In-person outreach to the general public reaches fewer people than other means, but it can yield impactful interactions, and may be perceived as having more value [Baucum, 2022]. As reported by one young attendee to in-person events: “one of the activities that most fostered my connection to the STEM world was attending in-person talks, conferences, and panel discussions […] ultimately leading me to pursue a degree in physics.”
@@
LVK members participate in science festivals, classroom visits, the creation of museum exhibits (such as interactive detector models [Cooper et al., 2021], or the Black Hole and Gravitational Wave circular exhibitions, lectures and panel discussions across Taiwan and Japan), visitor centers, and tours of the GW observatories. The face-to-face format allows staff to customize the experience to the visitor.
@@
LVK’s visitor centers merge artifacts from the detectors with interactive exhibits that put the visitor in the role of a scientist or engineer. The exhibits explore the underlying science and engineering of GW detectors, from the importance of pendula to concepts underpinning gravity. In-person and virtual tours of the detectors give a glimpse to the heart of operations. Over $330,000$ visitors have interacted with the LVK visitor centers; 2024 attendance is projected to reach $30,000$.
Science festivals and local outreach efforts around the world connect with classrooms and the general public. Outreach is performed by seasoned scientists, engineers, graduate students and undergraduates new to GW science. Such efforts positively impact the undergraduates who perform the outreach [Carpenter, 2015; Inverness Research Associates, 2014; Young & Katzman, 2023]. A 2023 survey revealed that $95\%$ of undergraduates said “their experience has changed the way they interact with their communities” and over half of LIGO Livingston program undergraduates said it “helped them decide to go into a STEM field or educational field” [Inverness Research Associates, 2023]. Undergraduates often note soft-skills development, one said “My experience as a LIGO docent has changed my life and how I interact with everyone. From conducting my presentations at school, teaching my kiddos, conducting business, or simply talking to others.”
2.6 Creating graphics
@@
Figure 4: The detector data for GW150914, and an artistic representation of its binary black hole source (credit: LIGO/National Science Foundation/Aurore Simmonet/Somona State University).
@@
Unlike electromagnetic observatories, GW data does not naturally lend itself to the stunning images that are commonplace throughout astronomy. Nevertheless, visualisations, infographics and cartoons [e.g., Kijbunchoo et al., 2016; Thompson et al., 2020] form key parts of our outreach activities, e.g., infographics displaying key facts accompany discoveries; simplified plots of results may be created for Science Summaries, and artistic representations of our sources may accompany press releases
@@
Videos and images are also produced from numerical-relativity simulations of sources and the GWs they emit. Since new simulations are often produced to better understand the physics of novel sources, fresh visualisations have accompanied discovery announcements.
@@
Images accompanying LVK discovery announcements have been selected as NASA Astronomy Pictures of the Day, e.g., to accompany the first detection, an artistic representation of a binary black hole coalescence (Figure 4) and a corresponding numerical-relativity simulation were featured (11 and 12 February 2016), and to accompany GW190521, an artistic representation of the GW emission from two spinning black holes was featured (8 September 2020).
2.7 Employing multisensory resources